OpenStudy (anonymous):

f(x)=3x-2 find 1) f(5x) , 2) 5f(x) for i) i got 15x-2 and for 2) i got 15x-10 is this correct?

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

f(5x) or f(5)?

OpenStudy (anonymous):

f(5x) or f(5)? 2Ash90

OpenStudy (anonymous):

the first one is f(5x)

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

then after you find that one out the second one we need to do is 5f(x)

OpenStudy (anonymous):

f(x)=3x-2 therefore f(5x)=3(5x)-2 =15x-2

OpenStudy (anonymous):

so far so good that's what i have for the first

OpenStudy (anonymous):

now for the second one i have 15x-10

OpenStudy (anonymous):

f(x)=3x-2 therefore 5f(x)=5(3x-2 ) =15x-10 thus 5f(x) =15x-10 @Ash90

OpenStudy (anonymous):

yes that is what i get also

OpenStudy (anonymous):

thanks for checking it for me

OpenStudy (anonymous):

@@Ash90 your both the answers are correct.

OpenStudy (anonymous):

thank you